OpenStudy (anonymous):

solve: |y-1| less than or equal to 8

OpenStudy (anonymous):

\[\le *\]

OpenStudy (anonymous):

\[-8 \le y - 1 \le 8\] add 1 to both sides..\[-7 \le y \le 9\]

OpenStudy (anonymous):

thannkkkss